OpenStudy (minecraftqueen):

Complete the standard form of the equation –4y = 3 – 2x. x + y = 3

OpenStudy (neonumbrella5115):

Do you know what standard form is?

OpenStudy (minecraftqueen):

Let's solve for x. −4y=3−2x Step 1: Flip the equation. −2x+3=−4y Step 2: Add -3 to both sides. −2x+3+−3=−4y+−3 −2x=−4y−3

OpenStudy (minecraftqueen):

Step 3: Divide both sides by -2. −2x −2 = −4y−3 −2

OpenStudy (minecraftqueen):

@Mehek14

OpenStudy (minecraftqueen):

@DanJS

OpenStudy (minecraftqueen):

@nincompoop